# Linking Process to Outcome: Conditional Reward Modeling for LLM Reasoning

### v1

@@ Abstract
Process Reward Models (PRMs) have emerged as a promising approach to enhance the reasoning capabilities of large language models (LLMs) by guiding their step-by-step reasoning toward a final answer. However, existing PRMs either treat each reasoning step in isolation, failing to capture inter-step dependencies, or struggle to align process rewards with the final outcome. Consequently, the reward signal fails to respect temporal causality in sequential reasoning and faces ambiguous credit assignment. These limitations make downstream models vulnerable to reward hacking and lead to suboptimal performance. In this work, we propose Conditional Reward Modeling (CRM) that frames LLM reasoning as a temporal process leading to a correct answer. The reward of each reasoning step is not only conditioned on the preceding steps but also explicitly linked to the final outcome of the reasoning trajectory. By enforcing conditional probability rules, our design captures the causal relationships among reasoning steps, with the link to the outcome allowing precise attribution of each intermediate step, thereby resolving credit assignment ambiguity. Further, through this consistent probabilistic modeling, the rewards produced by CRM enable more reliable cross-sample comparison. Experiments across Best-of-N sampling, beam search and reinforcement learning demonstrate that CRM consistently outperforms existing reward models, offering a principled framework for enhancing LLM reasoning. In particular, CRM is more robust to reward hacking and delivers stable downstream improvements without relying on verifiable rewards derived from ground truth.
@@ 1 Introduction
Recent advances in enhancing reasoning abilities have significantly improved the performance of large language models (LLMs) (Snell et al., 2025; Jaech et al., 2024), where models derive final answers through explicit step-by-step reasoning. Beyond prompt-based approaches (Wei et al., 2022; Diao et al., 2024; Fan et al., 2025), state-of-the-art systems like DeepSeek-R1 (Guo et al., 2025) have further advanced reasoning capacity through reinforcement learning (RL) with verifiable rewards.
However, verifiable rewards rely on checking model outputs against ground-truth labels, whose acquisition is costly and difficult to scale for general reasoning improvements. Reward models provide a promising alternative by extrapolating reward signals across general LLM reasoning processes. Broadly, these models fall into two categories: outcome reward models (ORMs) (Cobbe et al., 2021; Yu et al., 2024), which provide feedback at the final step; and process reward models (PRMs) (Wang et al., 2024; Li & Li, 2025; Yuan et al., 2025), which provide finer-grained rewards at the level of individual reasoning steps or even tokens.
Despite offering nuanced signals for reasoning processes, existing PRMs face several limitations. (i) Isolated step modeling: As shown in Figure 1(a), most PRMs (Lightman et al., 2023; Wang et al., 2024; Luo et al., 2024; Shao et al., 2024) assess each reasoning step in isolation, neglecting the intrinsic sequential dependencies of reasoning. (ii) Limited outcome awareness: While some methods (Yu et al., 2024; Li & Li, 2025; Yuan et al., 2025) attempt to mitigate isolated step rewards, they often fail to effectively link step-wise rewards with the outcome. For example, PQM (Li & Li, 2025), illustrated in Figure 1(a), relies on relative comparisons between neighboring steps (e.g., greater or smaller reward), but lacks explicit modeling of the final outcome, which may result in biased process rewards. IPRM (Yuan et al., 2025) parameterizes the outcome reward as the logarithmic sum of process rewards, yet it fails to capture how a specific step relate to the final outcome and lacks nuanced modeling of inter-step dependencies. This leads to ambiguous credit assignment from the final outcome back to the intermediate reasoning steps. As a result, existing methods are prone to reward hacking, where rewards continue to increase while the actual task accuracy declines, which we have observed in our experiments and has been shown Figure 1(b).
@@ 2 Related Work
To enhance LLM reasoning, prior studies have explored test-time search (e.g., Best-of-N, beam search) with reward models to select promising solutions (Xie et al., 2023; Zhang et al., 2025a; Snell et al., 2025), and post-training via RL (Jaech et al., 2024; Guo et al., 2025). Recent work (Zeng et al., 2025; Guo et al., 2025) has investigated RL with verifiable rewards for reasoning tasks. However, these approaches heavily rely on ground-truth and thus do not easily scale. Other studies (Cui et al., 2025; Cheng et al., 2025) integrate reward models into RL to provide dense feedback, but reward hacking remains a major concern (Gao et al., 2024).
@@
Most existing studies (Lightman et al., 2023; Wang et al., 2024; Luo et al., 2024; Shao et al., 2024) treat PRM as step-level classification, which treats steps in isolation and ignores inter-step dependencies. Recent studies (Lu et al., 2024; Li & Li, 2025; Yuan et al., 2025) attempts to move beyond classification-based PRMs. Li & Li (2025) reframes PRM as a Q-value ranking problem to model the relationships between steps, but overlook explicit modeling of the final outcome, resulting in a gap between process rewards and the final outcome. Yuan et al. (2025) instead proposes a parameterized formulation of outcome, leveraging outcome labels to train the PRM, but the relationship between intermediate step rewards and the final outcome remains unclear, and the lack of step dependency modeling leads to ambiguous credit assignment.
@@ 3.1 Task Description
Given a question $x$, LLM $\pi$ generates a response $y=(a_{1},a_{2},\ldots,a_{T})$ where $a_{t}$ denotes the $t$-th reasoning step and $T$ is the total step number. Let $a_{\leq t}$ denotes the sequence of the first $t$ steps. Each response is assigned a binary label $l\in\{0,1\}$ indicating whether the final answer derived from it is correct.
@@ 3.2 CRM Modeling
PBRS guarantees that any policy optimal under the shaped reward $R^{\prime}(s_{t},a_{t},s_{t+1})$ is also optimal under the original reward $R(s_{t},a_{t})$. Thus, using the shaped reward guides the learning process more effectively without changing the fundamental goal of the task. The key to applying PBRS lies in selecting an appropriate potential function $\Phi(s_{t})$. A natural choice for the potential function is one that estimates the likelihood of eventually reaching a correct answer from the current state. In our context, this corresponds to $S(t)$. We define the potential as the logarithm of this probability.
@@
The above derivation yields a dense and accurate process reward, $r_{t}=\log(1-h(t))$, which serves as a credit assignment scheme grounded in our modeling. The probability that the reasoning process reaches the correct final answer, $S(T)$, satisfies $S(T)=\prod_{t=1}^{T}(1-h(t))=\prod_{t=1}^{T}e^{r_{t}}$. Through this decomposition, we explicitly link process rewards to the outcome, addressing the limitation of prior work (Yu et al., 2024; Li & Li, 2025; Yuan et al., 2025) that lacks explicit modeling of the relationship between process and outcome.
@@ 3.3 Training of Conditional Reward Model
This consistent probabilistic modeling and training ensure that CRM is grounded in clear probabilistic semantics: for example, the value of $S(t)$ at any step $t$ across different samples consistently carries the same probabilistic meaning, allowing them to be compared under the same scale. By contrast, previous approaches (Li & Li, 2025; Yuan et al., 2025) struggle to achieve accurate cross-sample comparison (see Appendix E for detailed analysis).
@@ 4.1 Experimental Settings
Reward Model Training. We train the reward model on the Math-Shepherd dataset (Wang et al., 2024), which integrates questions from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) and provides responses annotated with step-level process labels. Following prior research (Li & Li, 2025; Guan et al., 2025), we augment the pre-trained model with a value head. The reward models are trained using full parameter fine-tuning. Further details can be found in Appendix A.
Baselines and Evaluation. We compare our CRM against representative baselines, including ORM, vanilla PRM (Wang et al., 2024), PQM (Li & Li, 2025) and IPRM (Yuan et al., 2025). For fairness, all baselines are re-implemented within the same pipeline, backbone, and training data. We then leverage the rewards provided by the reward model to (i) select an optimal response in Best-of-N (Section 4.2), (ii) guide beam search (Section 4.3), and (iii) optimize LLM reasoning through RL (Section 4.4).
@@ 4.2 Best-of-N Sampling Experiments
As shown in Table 1, CRM consistently ranks at or near the top across both datasets and model families. For example, on MATH500 with Qwen2.5-3B-Instruct, CRM reaches 56.6% at NN=32, surpassing the strongest baseline by +1.4 and remains ahead across NN=16, 64, and 128. By conditioning each step’s reward on all preceding steps, CRM ensures that the trajectory score reflects the coherence of the entire reasoning chain rather than isolated fragments. This holistic formulation enables CRM to assess logical consistency across steps and more reliably distinguish correct trajectories from superficial ones, leading to stronger trajectory-level selection.
@@ 4.4 RL Optimization Experiments
We compare against PRM and PQM, and additionally include two stronger RL baselines that use dense rewards: Prime (Cui et al., 2025) and PURE (Cheng et al., 2025). Prime adopts online reward model updates, jointly optimizing the reward model and policy, but its reliance on verifiers prevents adaptation to ground-truth–free settings. PURE adopts min-form credit assignment by defining the value function as the minimum of future rewards. However, as stated in its original paper, it becomes prone to reward hacking without verifiable rewards. To comprehensively evaluate performance, we use six widely adopted benchmarks: AIME25 (MAA, 2025), AIME24 (MAA, 2024), AMC23 (MAA, 2023), MATH500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). We report the pass@1 accuracy under the zero-shot setting.
@@
CRM-based RL functions without verifiable rewards (VR Disabled). For fair comparison with VR-dependent methods, we additionally report results where rewards from CRM are combined with verifiable rewards (VR Enabled). The results are shown in Table 3. In the VR Disabled setting, CRM attains the best Pass@1 accuracy on the majority of benchmarks, significantly outperforming the baselines. For instance, on AIME24 it reaches 43.3%, exceeding PURE by +16.7. When augmented with VR, CRM + VR yields further gains and achieves the highest performance on most benchmarks. This suggests that the process rewards provided by CRM are complementary to ground-truth–based verifiable rewards, rather than redundant. Remarkably, even without VR, CRM delivers performance comparable to VR-enabled methods. The strength of CRM lies in its explicit linkage between process rewards and outcomes, which ensures that the contribution of each step is causally aligned with the final result. Since process rewards quantify how much each step advances or undermines the probability of reaching the correct final answer, rather than relying on local heuristics, this design enables precise credit assignment. As a result, the dense and reliable reward signal facilitates effective RL optimization and remains robust to reward hacking, as further analyzed in Section 4.5.1.
@@ 4.5 Extended Analysis
In this section, we conduct a more in-depth analysis of CRM by investigating the following research questions. RQ1: What does reward hacking manifest as, and how can it be mitigated? RQ2: Does CRM exhibit self-reflection behavior? RQ3: How efficient is CRM in utilizing supervision data?
@@ 4.5.3 Ablation Study
Ablation on loss for CRM training. Our modeling approach employs three losses (Eq. 11, Eq. 12, Eq. 13). Among them, $\mathcal{L}_{S}$ and $\mathcal{L}_{W}$ are supervised by the label $l\in\{0,1\}$ indicating whether the final answer is correct, with a one-to-one correspondence to samples. Removing these losses would reduce the number of training samples and undermine fair comparison. Therefore, we preserve $\mathcal{L}_{S}$ and $\mathcal{L}_{W}$ while conducting ablations on $\mathcal{L}_{z}$. The loss $\mathcal{L}_{z}$ directly encourages the model to identify the exact step at which a wrong state occurs. Specifically, we set the total amount of data in the dataset applicable to $\mathcal{L}_{z}$ as 100% and apply scaling, evaluating CRM with 0%, 25%, 50%, and 75% of the data under the Best-of-N sampling on the MATH500 dataset using Qwen2.5-3B-Instruct.
@@
Proportion of
data used for $\mathcal{L}_{z}$
@8 @16 @32 @64 @128
@@
Finding 3: CRM exhibits high data efficiency (RQ3). The ablation results in Table 4 show that even a very small proportion of data applicable to $\mathcal{L}_{z}$ leads to substantial performance gains: moving from 0% to just 10% supervision produces a large improvement in Best-of-N accuracy. Increasing the proportion of data yields additional benefits, but the improvements quickly saturate, with 50% already achieving near-optimal performance. This demonstrates that CRM requires only limited data for $\mathcal{L}_{z}$ to achieve strong results, highlighting its high data utilization efficiency.

### v2

# arXiv:2509.26578v2[cs.LG] 27 Feb 2026

@@ Abstract
Process Reward Models (PRMs) have emerged as a promising approach to enhance the reasoning capabilities of large language models (LLMs) by guiding their step-by-step reasoning toward a final answer. However, existing PRMs either treat each reasoning step in isolation, failing to capture inter-step dependencies, or struggle to align process rewards with the final outcome. Consequently, the reward signal fails to respect temporal causality in sequential reasoning and faces ambiguous credit assignment. These limitations make downstream models vulnerable to reward hacking and lead to suboptimal performance. In this work, we propose Conditional Reward Modeling (CRM) that frames LLM reasoning as a temporal process leading to a correct answer. The reward of each reasoning step is not only conditioned on the preceding steps but also explicitly linked to the final outcome of the reasoning trajectory. By enforcing conditional probability rules, our design captures the causal relationships among reasoning steps, with the link to the outcome allowing precise attribution of each intermediate step, thereby resolving credit assignment ambiguity. Further, through this consistent probabilistic modeling, the rewards produced by CRM enable more reliable cross-sample comparison. Experiments across Best-of-N sampling, beam search and reinforcement learning demonstrate that CRM consistently outperforms existing reward models, offering a principled framework for enhancing LLM reasoning. In particular, CRM is more robust to reward hacking and delivers stable downstream improvements without relying on verifiable rewards derived from ground truth. The project website is at https://foundation-model-research.github.io/CRM.
@@ 1 Introduction
Recent advances in enhancing reasoning abilities have greatly improved the performance of large language models (LLMs) (Snell et al., 2025; Jaech et al., 2024), where models derive final answers through explicit step-by-step reasoning. Beyond prompt-based approaches (Wei et al., 2022; Diao et al., 2024; Fan et al., 2025), state-of-the-art systems like DeepSeek-R1 (Guo et al., 2025a) have further advanced reasoning capacity through reinforcement learning (RL) with verifiable rewards.
However, verifiable rewards rely on checking model outputs against ground-truth labels, whose acquisition is costly and difficult to scale for general reasoning improvements. Reward models provide a promising alternative by extrapolating reward signals across general LLM reasoning processes. Broadly, these models fall into two categories: outcome reward models (ORMs) (Cobbe et al., 2021; Yu et al., 2024), which provide feedback at the final step; and process reward models (PRMs) (Wang et al., 2024; Li and Li, 2025; Yuan et al., 2025), which provide finer-grained rewards at the level of individual reasoning steps or even tokens.
Despite offering nuanced signals for reasoning processes, existing PRMs face several limitations. (i) Isolated step modeling: As shown in Figure 1(a), most PRMs (Lightman et al., 2023; Wang et al., 2024; Luo et al., 2024; Shao et al., 2024) assess each reasoning step in isolation, neglecting the intrinsic sequential dependencies of reasoning. (ii) Limited outcome awareness: While some methods (Yu et al., 2024; Li and Li, 2025; Yuan et al., 2025) attempt to mitigate isolated step rewards, they often fail to effectively link step-wise rewards with the outcome. For example, PQM (Li and Li, 2025), illustrated in Figure 1(a), relies on relative comparisons between neighboring steps (e.g., greater or smaller reward), but lacks explicit modeling of the final outcome, which may result in biased process rewards. IPRM (Yuan et al., 2025) parameterizes the outcome reward as the logarithmic sum of process rewards, yet it fails to capture how a specific step relates to the final outcome and lacks nuanced modeling of inter-step dependencies. This leads to ambiguous credit assignment from the final outcome back to the intermediate reasoning steps. As a result, existing methods are prone to reward hacking, where rewards continue to increase while the actual task accuracy declines, which we have observed in our experiments and has been shown in Figure 1(b).
@@ 2 Related Work
To enhance LLM reasoning, prior studies have explored test-time search (e.g., Best-of-N, beam search) with reward models to select promising solutions (Xie et al., 2023; Zhang et al., 2025a; Snell et al., 2025), and post-training via RL (Jaech et al., 2024; Guo et al., 2025a). Recent work (Zeng et al., 2025b; Guo et al., 2025a) has investigated RL with verifiable rewards for reasoning tasks. However, these approaches heavily rely on ground-truth and thus do not easily scale. Other studies (Cui et al., 2025; Cheng et al., 2025) integrate reward models into RL to provide dense feedback, but reward hacking remains a major concern (Gao et al., 2024).
@@
Most existing studies (Lightman et al., 2023; Wang et al., 2024; Luo et al., 2024; Shao et al., 2024) treat PRM as step-level classification, which treats steps in isolation and ignores inter-step dependencies. Recent studies (Lu et al., 2024; Li and Li, 2025; Yuan et al., 2025) attempt to move beyond classification-based PRMs. Li and Li (2025) reframes PRM as a Q-value ranking problem to model the relationships between steps, but overlooks explicit modeling of the final outcome, resulting in a gap between process rewards and the final outcome. Yuan et al. (2025) instead proposes a parameterized formulation of outcome, leveraging outcome labels to train the PRM, but the relationship between intermediate step rewards and the final outcome remains unclear, and the lack of step dependency modeling leads to ambiguous credit assignment. Another line of work (Chen et al., 2025; Guo et al., 2025b; She et al., 2025; Zhao et al., 2025) integrates reasoning capabilities into reward modeling, such as R-PRM (She et al., 2025) and GenPRM (Zhao et al., 2025), which evaluate reasoning processes with detailed thinking. However, these approaches often involve complex training procedures and incur high inference costs.
@@ 3.1 Task Description
Given a question $x$, LLM $\pi$ generates a response $y=(a_{1},a_{2},\ldots,a_{T})$ where $a_{t}$ denotes the $t$-th reasoning step and $T$ is the total step number. Let $a_{\leq t}$ denote the sequence of the first $t$ steps. Each response is assigned a binary label $l\in\{0,1\}$ indicating whether the final answer derived from it is correct.
@@ 3.2 CRM Modeling
PBRS guarantees that any policy optimal under the shaped reward $R^{\prime}(s_{t},a_{t},s_{t+1})$ is also optimal under the original reward $R(s_{t},a_{t})$. Thus, using the shaped reward guides the learning process more effectively without changing the fundamental goal of the task. The key to applying PBRS lies in selecting an appropriate potential function $\Phi(s_{t})$ which is effective when it encodes goal-directed progress (Lo et al., 2024; Müller and Kudenko, 2025; Mindom et al., 2021; Wiewiora et al., 2003). We define the potential function as $\log S(t)$, which denotes the log-likelihood of eventually reaching a correct answer from the current state, thereby explicitly capturing the progress toward the final correct answer. Consequently, substituting this definition of $\log S(t)$ into the potential function yields the following explicit form:
@@
The above derivation yields a dense and accurate process reward, $r_{t}=\log(1-h(t))$, which serves as a credit assignment scheme grounded in our modeling. The probability that the reasoning process reaches the correct final answer, $S(T)$, satisfies $S(T)=\prod_{t=1}^{T}(1-h(t))=\prod_{t=1}^{T}e^{r_{t}}$. Through this decomposition, we explicitly link process rewards to the outcome, addressing the limitation of prior work (Yu et al., 2024; Li and Li, 2025; Yuan et al., 2025) that lacks explicit modeling of the relationship between process and outcome.
@@ 3.3 Training of Conditional Reward Model
This consistent probabilistic modeling and training ensure that CRM is grounded in clear probabilistic semantics: for example, the value of $S(t)$ at any step $t$ across different samples consistently carries the same probabilistic meaning, allowing them to be compared under the same scale. By contrast, previous approaches (Li and Li, 2025; Yuan et al., 2025) struggle to achieve accurate cross-sample comparison (see Appendix E for detailed analysis).
@@ 4.1 Experimental Settings
Reward Model Training. We train the reward model on the Math-Shepherd dataset (Wang et al., 2024), which integrates questions from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) and provides responses annotated with step-level process labels. Following prior research (Li and Li, 2025; Guan et al., 2025), we augment the pre-trained model with a value head. The reward models are trained using full parameter fine-tuning. Further details can be found in Appendix A.
Baselines and Evaluation. We compare our CRM against representative baselines, including ORM, vanilla PRM (Wang et al., 2024), PQM (Li and Li, 2025) and IPRM (Yuan et al., 2025). For fairness, all baselines are re-implemented within the same pipeline, backbone, and training data. We then leverage the rewards provided by the reward model to (i) select an optimal response in Best-of-N (Section 4.2), (ii) guide beam search (Section 4.3), and (iii) optimize LLM reasoning through RL (Section 4.4).
@@ 4.2 Best-of-N Sampling Experiments
As shown in Table 1, CRM consistently ranks at or near the top across both datasets and model families. For example, on MATH500 with Qwen2.5-3B-Instruct, CRM reaches 56.6% at NN=32, surpassing the strongest baseline by +1.4 and remaining ahead across NN=16, 64, and 128. By conditioning each step’s reward on all preceding steps, CRM ensures that the trajectory score reflects the coherence of the entire reasoning chain rather than isolated fragments. This holistic formulation enables CRM to assess logical consistency across steps and more reliably distinguish correct trajectories from superficial ones, leading to stronger trajectory-level selection. A more in-depth analysis of the performance characteristics of BoN is provided in Appendix I.
@@ 4.4 RL Optimization Experiments
We compare against PRM and PQM, and additionally include two stronger RL baselines that use dense rewards: Prime (Cui et al., 2025) and PURE (Cheng et al., 2025). Prime adopts online reward model updates, jointly optimizing the reward model and policy, but its reliance on verifiers prevents adaptation to ground-truth–free settings. PURE adopts min-form credit assignment by defining the value function as the minimum of future rewards. However, as stated in its original paper, it becomes prone to reward hacking without verifiable rewards. In addition, we also add the VR baseline, where the policy is trained using only final-answer verifiable rewards. To comprehensively evaluate performance, we use six widely adopted benchmarks: AIME25 (MAA, 2025), AIME24 (MAA, 2024), AMC23 (MAA, 2023), MATH500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). We report the pass@1 accuracy under the zero-shot setting.
@@
CRM-based RL functions without verifiable rewards (VR Disabled). For fair comparison with VR-dependent methods, we additionally report results where rewards from CRM are combined with verifiable rewards (VR Enabled). The results are shown in Table 3. In the VR Disabled setting, CRM attains the best Pass@1 accuracy on the majority of benchmarks, significantly outperforming the baselines. For instance, on AIME24 it reaches 43.3%, exceeding PURE by +16.7. CRM consistently outperforms the VR baseline. When augmented with VR, CRM + VR yields further gains and achieves the highest performance on most benchmarks. This suggests that the process rewards provided by CRM are complementary to ground-truth–based verifiable rewards, rather than redundant. Remarkably, even without VR, CRM delivers performance comparable to VR-enabled methods. The strength of CRM lies in its explicit linkage between process rewards and outcomes, which ensures that the contribution of each step is causally aligned with the final result. Since process rewards quantify how much each step advances or undermines the probability of reaching the correct final answer, rather than relying on local heuristics, this design enables precise credit assignment. As a result, the dense and reliable reward signal facilitates effective RL optimization and remains robust to reward hacking, as further analyzed in Section 4.5.1.
@@ 4.5 Extended Analysis
In this section, we conduct a more in-depth analysis of CRM by investigating the following research questions. RQ1: What does reward hacking manifest as, and how can it be mitigated? RQ2: Does CRM exhibit self-reflection behavior? RQ3: How efficient is CRM in utilizing supervision data? RQ4: Can CRM be applied to domains beyond mathematics?
@@ 4.5.3 Ablation Study
Proportion of data used for $\mathcal{L}_{z}$ @8 @16 @32 @64 @128
@@
Ablation on loss for CRM training. Our modeling approach employs three losses (Eq. 11, Eq. 12, Eq. 13). Among them, $\mathcal{L}_{S}$ and $\mathcal{L}_{W}$ are supervised by the label $l\in\{0,1\}$ indicating whether the final answer is correct, with a one-to-one correspondence to samples. Removing these losses would reduce the number of training samples and undermine fair comparison. Therefore, we preserve $\mathcal{L}_{S}$ and $\mathcal{L}_{W}$ while conducting ablations on $\mathcal{L}_{z}$. The loss $\mathcal{L}_{z}$ directly encourages the model to identify the exact step at which a wrong state occurs. Specifically, we set the total amount of data in the dataset applicable to $\mathcal{L}_{z}$ as 100% and apply scaling, evaluating CRM with 0%, 10%, 25%, and 50% of the data under the Best-of-N sampling on the MATH500 dataset using Qwen2.5-3B-Instruct.
Finding 3: CRM exhibits high data efficiency (RQ3). The ablation results in Table 4 show that even a very small proportion of data applicable to $\mathcal{L}_{z}$ leads to substantial performance gains: moving from 0% to just 10% supervision produces a large improvement in Best-of-N accuracy. Increasing the proportion of data yields additional benefits, but the improvements quickly saturate, with 50% already achieving near-optimal performance. This demonstrates that CRM requires only limited data for $\mathcal{L}_{z}$ to achieve strong results, highlighting its high data utilization efficiency. A more detailed analysis of the reasons behind this efficiency, as well as comparisons with the baseline, can be found in the Appendix  J.
4.5.4 Validation on More Domains
To further validate the effectiveness of our CRM, we extend our evaluation to additional domains beyond mathematics. We leverage the multi-domain dataset MMLU-Pro-CoT-Train from VersaPRM (Zeng et al., 2025a) as it provides reasoning processes with step-wise labels across a wide range of domains. We train the reward models including CRM and the baselines on this dataset using Qwen2.5-3B-Instruct as backbone and subsequently assess their performance using Best-of-N sampling on five distinct domains including biology, business, health, history, and physics.
Finding 4: CRM can be extended to other domains and demonstrates strong performance. (RQ4). As shown in Table 5, CRM outperforms the baselines in almost all tested domains. This demonstrates that CRM’s validity is not limited to mathematical tasks and can be applied to a broader range of domains.
Table 5: Best-of-N accuracy comparison across different domains (MMLU-Pro-CoT-Eval).
Methods Biology Business Health History Physics
@16 @32 @64 @16 @32 @64 @16 @32 @64 @16 @32 @64 @16 @32 @64
PRM 80.0 79.2 79.2 64.8 64.8 66.2 60.7 57.1 57.1 48.0 48.7 48.7 61.0 58.9 60.3
PQM 80.8 80.0 80.8 64.1 63.5 65.5 62.1 61.4 61.4 50.0 51.3 49.3 61.6 58.2 62.3
[…]